in a rose flower (angiosperm) the unfertilized egg has 18 chromosomes. later the egg is fertilized by a sperm.
how many chromosomes does the embryo have?
how many chromosomes does the endosperm have?
how many chromosomes does the integument have?
Answer #1

1. In angiosperms, one sperm cell fertilizes an egg and forms an embryo and other sperm cell fuses with polar nuclei in the embryo sac and forms endosperm. This is called double fertilization. An embryo is formed when an egg (n) is fertilized by sperm (n), therefore embryo is diploid (2n). It has two sets of chromosomes.

n = one set of chromosomes = 18

number of chromosomes in embryo = 2n = 2* 18 = 36

Embryo (diploid) has 36 chromosomes

2. Endosperms are formed as a result of double fertilization in angiosperms. When sperm (n) fuses with polar nuclei (2n), the endosperm is formed. Hence endosperms in angiosperms are triploid (3n). It has 3 sets of chromosomes

number of chromosomes in endosperm = 3n = 3 * 18 = 54

Endosperm (triploid) has 54 chromosomes

3. The integument is an outer covering of an ovule. In angiosperms, there are two integuments. Integument of the ovule of angiosperm is diploid (2n). It has two sets of chromosomes.

number of chromosomes in integument = 2n = 2* 18 = 36

Integument (diploid) has 36 chromosomes